OpenStudy (anonymous):

calc help! find the extrema on the closed interval. g(x)=2x+5cosx

zepdrix (zepdrix):

What interval? c:

OpenStudy (anonymous):

g(x)=2x+5cosx once derived...its... 2-5sinx

OpenStudy (anonymous):

o wait ! [0, 2pi]

OpenStudy (anonymous):

sorry

zepdrix (zepdrix):

\[g'(x)=2-5\sin x\]\[0=2-5\sin x\]\[\sin x=\frac{ 2 }{ 5 }\]\[\arcsin \frac{ 2 }{ 5 }=x\] So this is how we solve for critical points yes? :o

zepdrix (zepdrix):

Which is not a nice special angle :[ grr

OpenStudy (anonymous):

idk about the arcsin?????

zepdrix (zepdrix):